Find the equation of the line through the point (6,6) and parallel to the point y=6-(x/2). Express the equation of the line in slope-intercept form, y=mx+b.
My first step was to put the (6,6) into y=mx+b
6=-(1/2)(6)+b -> 6=-(6/2)+b -> b=(18/2)? so then y=-(1/2)x+(18/2)
If someone wouldn't mind going over the work I would greatly appreciate it :)

Found 3 solutions by Alan3354, ikleyn, math_tutor2020:
Answer by Alan3354(69443)

Find the equation of the line through the point (6,6) and parallel to the point y=6-(x/2). Express the equation of the line in slope-intercept form, y=mx+b.
My first step was to put the (6,6) into y=mx+b
6=-(1/2)(6)+b -> 6=-(6/2)+b -> b=(18/2)? so then y=-(1/2)x+(18/2)
If someone wouldn't mind going over the work I would greatly appreciate it :)
-------------------------
Find the slope of y = 6-(x/2)
In slope-intercept form: y = (-1/2)x + 6
The slope is -1/2
Parallel lines have the same slope.
Use y-y1 = m*(x-x1) where m is the slope and (x1,y1) is the point
y-6 = (-1/2)*(x-6)
y = (-1/2)x + 9 --- same as your answer.

Answer by ikleyn(52781)

.
At this web-site,  there is a group of lessons
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
that are closely related to your problem.

They cover the entire subject with all inter-related problems in one way.

You have a happy opportunity to learn the subject from these lessons.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.



Answer by math_tutor2020(3817)


It's best to reduce 18/2 to 9.

Therefore, the answer would be y = (-1/2)x+9

As a check, plug the coordinates of (6,6) into that equation.
y = (-1/2)x+9
6 = (-1/2)*6+9
6 = -3+9
6 = 6
Both sides result in the same value, so we get a true equation.
This confirms (6,6) is on the line y = (-1/2)x+9

Furthermore, the two lines y = 6-(x/2) and y = (-1/2)x+9 are parallel since they have the same slope -1/2.

y = 6-(x/2) in green
y = (-1/2)x+9 in blue

I recommend using a tool like Desmos or GeoGebra to graph.